FEL Upgrade Project Weekly Brief

January 21-25, 2002




This week we made significant progress with the installation of new hardware in the high voltage power supply for the Upgrade injector RF power supply and we received the first 20 of the Upgrade quadrupole magnets.


FEL Project management along with the majority of the management of Jefferson Lab were consumed this week by a critical DOE review of the laboratory’s operational budget and budget projections for the next 4 years. This review will be used by DOE’ Office of Nuclear Physics to benchmark the costs for CEBAF operations, associated operational support costs, and the necessary costs to keep the lab’s core competencies (i.e., SRF and source technology) viable. Based on the preliminary report given by the review committee, the lab was viewed as a well run and lean operation. In addition, constant effort budgets for the next 4 years would lead to degradations in machine performance and inadequate support of core technologies. Hopefully,

this analysis will help DOE enlist Congressional and Administration support for additional funds.

A draft revision of our current DOE/ONR MOA for the FEL activities was prepared by our ONR contract manager and directed to us and the DOE site office for comments. The current MOA is being revised to encompass new activities that will be funded in FY02 and FY03 including upgrade commissioning, completion of the third cryomodule, design and construction of a new wiggler optimized for 1 micron operation and supplemental hardware projects funded by the JTO. (Note: two JTO projects involving cryogenic mirror design and drive laser design were recently awarded to the FEL project).

We are pleased to congratulate Old Dominion University, one of our Applied Research Center partners, for successfully soliciting sufficient matching funds from partner organizations to meet the final requirements for a NSF Center for Industry-University Collaborative Research on laser-material interactions. ODU expects to hear an official announcement of the award for the center from NSF next month. We are proud of our colleagues at ODU for this achievement and believe the Center activities will be very useful to FEL users.





WBS 3 (Beam Physics):

Analysis of magnet modeling results continues. 2F and 5F diagnostics configurations were evaluated and appear acceptable.

WBS 4 (Injector):

The gun chamber is having the upstream and downstream tubes machined. Parts for the rear portion of the shield operating mechanism were completed by the shop, some welding still required. The front end shield assembly was baked to 150C for water removal and still operates fine under these vacuum conditions. The drawings of the extractor cart for the ceramic stack need modifications to enable use if we go to the 10 MV/m configuration. Design drawings for the inline resistor are being worked on.

Gun HVPS - The HVPS Tank was reviewed again this week and is undergoing some modifications. The drawings should be ready for signatures next week. All parts have been removed form the tank and stored in the Faraday room in the Test Lab building. The tank is ready for removal and storage in the tent behind the FEL building.

Gun Resistors - The resistive, terminal, and coating materials for the Conditioning Resistor have been received. The wire cloth for screening is ordered and is due early next week. The experiments will start shortly thereafter.

WBS 6 (RF):

Quarter HVPS's - The mechanical installation is about 80% complete with electrical wiring to start on Monday. The first cut at the schematic is being completed by Drafting. A working 50 kW klystron will be used as a load to checkout the HVPS to 28 kV.

Quarter Klystrons - No progress for the past 2 weeks in modifying the LCW manifolds by the Machine Shop. They have promised delivery by next Thursday. The 2.5" taps in the LCW lines are to be installed in February.

WBS 8 (Instrumentation):

Cables have been removed from the cable trays for the necessary modifications in the new optical cavity regions.

Nomenclature and elements for vacuum, and diagnostic components has been reviewed and drawings updated.

Six new three inch shielded beamviewers have been assembled for turnover to the vacuum group for girder assembly.

Drawings are being updated for the new stepper motor control chassis and beamviewer camera light rings. Assembly of these items should begin soon.

Eleven more ion pump power supplies have been received from the assembler and will be inspected and tested next week.



WBS 9 (Transport):


Optical Chicane Dipoles (GW)

o I visited Magnet Enterprises International in Oakland CA. They successfully wound the new first coil referred to last week, bent the ears up to form the saddle shape and are working on leads and ground wrap. Dimensional checks show no problem with the coil fitting in the gap, or being larger than the calculated keystone expansion of the bend-up of the ears. Their potting mold is ready for potting this coil, they will utilize the temperature data from the prototype coil combined with additional real time temperature probe readings to maintain the required temperature cycle for potting.

Cores parts are in machining by vendors experienced at grinding and machining magnets made of annealed 1006 steel. (Several SLAC magnet cores were just being completed at F &F Grinding during my visit.) All other parts are accounted for by being purchased or by material on order for fabricate in-house.

Injector Dipoles (DU/DV)

o Procurement is continuing via electronic methods. We have granted a three day extension to the bid time because of several requests by vendors.

o DULY research, in checking their magnetic model, found that the Small Injector Dipole (GV) field integral quality is a strong function of the coil saddle ear’s distance from the inner surface of the field clamp’s horizontal plate. (Not so for the GU larger dipole) It is difficult to maintain tight tolerance on the position of this coil surface. We are being prudent, asking for a detailed sensitivity analysis. They will alter the drawings to add a shim set on the inside of the field clamp. Magnet Measurement Folks will be able to alter the coil to field clamp distance to tweak the magnet into magnetic tolerance by adjusting the quantity of shims.

o All drawing revisions will have to be done by the time we place of the order for these magnets.

Arc Dipoles (GY, GX, GQ)

o Advanced Energy Systems (AES) analyzed the field integral information on the Reverse Bend Magnet (GQ) at larger field position and at different levels of excitation. David Douglas is reviewing their findings.

o I visited Wang NMR of CA who are manufacturing the 180° Dipole coils. They are starting to insulate the conductor and design the winding form and potting mold. I verified that their anticipated winding, potting and curing methods were realistic for this odd shape coil.


3 inch quad (QX)

o Twenty QX Quadrupoles arrived. We tried to qualify the First Article Quadrupole. However, the magnet measurement stand didn't automatically center the quad on the fiducials. We are investigating the cause.

o New England Technicoil has used up all the coils for assembling quadrupoles. They have shipped several additional batches of ten magnets.

Trim Quad (QT)

o Tom Hiatt visited Milhous Control of Virginia to have a discussion of critical parameters before actual production begins.

Sextupole (SF)

o DULY has created efficient field clamps on the sextupole’s magnetic model and met specification at course mesh. The final design at fine mesh was postponed because of the more urgent need for a re-look at the GV magnetic model.

Octupole (OT)

o Work on this magnet is on furlough until the sextupole is designed.

Corrector Dipoles (DB, DJ, GC BH)

o No news on qualification of the new sets of Standard Corrector Dipoles (DB/DJ). QX Quadrupole has the priority on the test stand.

o DULY Research successfully magnetically modeled the intended simple new vertical corrector for the Arcs (GC).

o Jeff Dale of the Shop started the necessary planning to make the single set of bakable Haimson style correctors that is needed for the new gun.

Beam Line and Vacuum

o AES continues drawing up the chambers through out the arcs and is completing finite element analysis of the most problematic chamber, the chamber at the GX in the optical cavity. It mandates a 5/16 inch thick plate.

o Design work on The Optical Chicane region chamber, stands and diagnostics continued.

o David Douglas reviewed and validated his latest DIMAD revision of the 6 quad telescopes just before the first arc and after the Second Arc. (2F & 5F Regions) that were changed because of UV concerns. He is reviewing the shoe-horning of the diagnostics into the regions.

WBS 11 (Optics):

The revised drawing package for the optical cavity assemblies is being checked. The mirror mount for the ultraviewer was revised to make it less prone to movement during an undamped excursion; these drawings are in review as well. The optical cavity stepper motors test fixture is assembled and ready for use. More tests with Ga:In eutectic material show it wets molybdenum, we are observing the mount to see if there is any reaction. The Thompson rails for the optical cavity assemblies were received this week. We made more progress in understanding bubble formation and bonding uniformity, as a function of application, for the HR deformable mirror assemblies. Answers to questions raised in last month's review of the OCMMS design will be distributed for comment; assuming all questions are answered satisfactorily, we will finish the drawing package. Tests of the new position sensitive detector (PSD) began in the OCMMS test stand. Modeling of a scaled up high power (50kW) laser beam dump, based on our previous design, has recommenced. It will be compared in projected performance and cost with a quote from an outside vendor.

At the Photonics West meeting this week, we met with vendors of drive laser components (oscillator and amplifiers). A great deal of information was obtained, and will help in solidifying our design.

Operations/Other Activities:

Lights were replaced in the Drive Laser Enclosure. Plans were made for moving the optical tables, used in the vault, upstairs to the User Labs.